OpenStudy (anonymous):

Need statistics help again. When a truckload of apples arrives at a packing plant, a random sample of 200 is selected and examines for bruises, discolorations etc. The whole truckload will be rejected if more than 6% of the sample is unsatisfactory. Suppose that in fact 9% of the apples on the truck do not meet the desired standard. What is the probability that the shipment will be accepted anyway?

OpenStudy (kropot72):

This can be solved by using the normal approximation to the binomial distibution. The mean is np=200 * 0.09 = 18 The standard deviation is \[\sqrt{np(1-p)}=\sqrt{200\times 0.09\times 0.91}=4.04722\] Now we need to find the z-score for 12 which corresponds to 6% of the apples being unsatisfactory. \[z=\frac{X-\mu}{\sigma}=\frac{12-18}{4.04722}\] The cumulative probability for the z-score is found by using a standard normal distribution table.

OpenStudy (anonymous):

Thank you very much!

OpenStudy (kropot72):

You're welcome :)